Badlands, Inc. manufactures a household fan that sells for $25 per unit. All sales are on account, with 30 percent of sales the month of sale and 70 percent collected in the following month. The data that follow were extracted from the companys accounting in . . Badlands maintains a minimum cash balance of $21,000. Total payments in January 20x1 are budgeted at $205,000. A schedule of cash for January and February of 20x1 revealed the following receipts for the period: $105,000 91,00a $144,eee From February sales March 20x1 sales are expected to total 7000 units Finished-goods inventories are maintained at 30 percent of the following months sales . The 31, 20x0, balance sheet revealed the following selected figures: cash, $24,000 accounts receivable, $105,000 and finished goods, $24.600. References Required: the number of units that Badlands sold in December 20x0 the sales revenue for March 20x1 e the total sales revenue to be reported on Badlands budgeted income statement for the first quarter of 20x1 4. Determine the accounts receivable balance to be reported on the March 31, 20x1, budgeted balance sheet 5. Calculate the number of units in the December 31. 20xO, finished-goods inventory e the number of units of finished goods to be manufactured in January 20xt, 7. Calculate the financing required in January, if any, to maintain the firms minimum cash ba 1. D inventory 7 Financing required in January
Answer #1
1) December sales =
Accounts receivable /(1-0.30) * Selling price per unit
105000/0.70*25 = 6000 units
2) March sales revenue= units sold in march * selling price per unit =
7000 units * $25 per unit = 175000
3) First quarter sales revenue=
jan sales + feb sales + mar sales=
(91000+144000)+(70500/30%)+175000 = 645000
4) March 31 accounts receivable balance =
175000 * 70% = $122500
5) December 31 finished goods inventory =
December closing inventory / Unit sales per unit
$24600 / $25 per unit = 984 units
6) Finished goods to be manufactured in January =
(Jan sales in units + closing inventory - opening inventory)=
(235000/25) + (70500*30%/30%*25) - 984 =
9400 + 2820 - 984 = 11236 units
7) Financing required in Jan =
OB of cash 24000
add collections 196000
less payments -205000
Excess of cash over payment 15000
less: minimum cash reqd. 21000
Financing required $6,000